OpenStudy (anonymous):

Is my answer correct guys???? THANKS! Solve. 3x2 - 7x - 2 = 0 answer below:

OpenStudy (anonymous):

x = 7 + square root 73/ 6

OpenStudy (anonymous):

Are you sure the signs are right in the original equation?

OpenStudy (anonymous):

3x^2 - 7x - 2 = 0

jimthompson5910 (jim_thompson5910):

if you mean to say \[\Large x = \frac{7\pm\sqrt{73}}{6}\] then you are correct

OpenStudy (anonymous):

thank you so much!!!!!!!!! :)

jimthompson5910 (jim_thompson5910):

np